OpenStudy (anonymous):

What is the following number in decimal form? 8.61 × 10 –8 A. 0.000000861 B. 0.0000000861 C. 8,610,000 D. 861,000,000

OpenStudy (igreen):

First you should simplify \(10^{-8}\).

OpenStudy (anonymous):

i think i got it

OpenStudy (igreen):

\(10^{-8} \rightarrow \dfrac{1}{10^8}\)

OpenStudy (anonymous):

ummm....i think the answer is D. Am i correct?

OpenStudy (anonymous):

because....

OpenStudy (anonymous):

let me explain

OpenStudy (igreen):

No, actually C and D are wrong..because our answer will be a decimal.

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

10^-8 = -10 x -10 x -10 x -10 x -10 x -10 x -10 x-10

OpenStudy (anonymous):

which gives me 100,000,000

OpenStudy (anonymous):

then 8.61 x 100,000,000 gives me 861,000,000

OpenStudy (anonymous):

or C thats the ones i am thinking because it isn't 100,000,000 x 8.61 it is 8.61 x 100,000,000

OpenStudy (anonymous):

and i get C

OpenStudy (anonymous):

Confused plz help

OpenStudy (anonymous):

i only think D is the answer but not so sure anymore

OpenStudy (anonymous):

what do you think and why

myininaya (myininaya):

\[10^2=100 \\ \text{ where as } \\ 10^{-2}=0.01\] so your 10^(-8) will be a decimal number as suggested above

OpenStudy (mathstudent55):

When you multiply a number by an integer power of ten, look at the exponent of 10. If the exponent is positive, move the decimal point of the number that number of places to the right. If the exponent is negative, move the decimal point that number of places to the left.

OpenStudy (anonymous):

How

OpenStudy (anonymous):

oh

OpenStudy (mathstudent55):

Examples: 5.2 * 10^5 = 520,000 3.6 * 10^-4 = 0.00036

OpenStudy (anonymous):

B?

OpenStudy (mathstudent55):